OpenStudy (bananas1234):

H(x) = 16 - x, when x = -4

OpenStudy (mathmale):

Here's what I'd recommend you do (to avoid problems with negative signs): 1. Re-write H(x) = 16 - x as H(x) = 16 - ( x ) 2. Substitute -4 for x (inside those parentheses) 3. Show your work and your final result.

OpenStudy (mathmale):

It'd be wise to write your result as H(-4) = 16 - (-4). H(-4) reads as "the value of function H(x) when x=-4."

OpenStudy (vuriffy):

A double negative configures into a positive value.

OpenStudy (vuriffy):

Understand?